I would like some help with this problem:
The admission to the Coney Island Amusement Park is $1.80 for children, and $5.00 for adults. On Saturday, 2,500 people went to the amusement park and $6,050 was collected. How many children and adults attended the amusement park?

Answer by Alan3354(69443)

The admission to the Coney Island Amusement Park is $1.80 for children, and $5.00 for adults. On Saturday, 2,500 people went to the amusement park and $6,050 was collected. How many children and adults attended the amusement park?
--------------
c + a = 2500 (2500 children + adults)
180c + 500a = 605000 (total admission charge)
a = 2500 - c (from eqn 1)
Sub into eqn 2
180c + 500(2500-c) = 605000
180c + 1250000-500c = 605000
-320c = - 645000
c = 2015.625 children
a = 484.375
-------------
I would expect integers, I suspect a typo.